Ticket Number: HEA-NOC/20050409-2 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20050409 12:43 UTC+1 Problem Start: 20050409 12:13 UTC+1
Ticket Closed: 20050411 11:30 UTC+1 Problem End: 20050409 12:47 UTC+1
Site/line: Digital Depot
The bluenet link from TCD to the Digital Depot has gone down. Cannot connec
to the Digital Depot Cisco 3550.
Client RIFTFS has lost connectivity to HEAnet.
20050409 12:26 EK Contacted the digital Depot NOC. No maintenance is
planned for the Depot this weekend. Digital Depot will see if local hands
can be arranged.
Contacted Dublin City Council Traffic department, all their fibre links
are working. Traffic has no record of any outages.
20050409 13:14 EK link came back up at 12:47. Digital depot noc will
advise on reason for outage.
20050411 11:24 EK Digital Depot NOC have informed us that unautorized
access to the Digital Depot comms room on Saturday by TKO staff, resulted
in the outage. The outage was caused by our fibre in the Depot being
disconnected from our equipment. TKO access to the comms room has been
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20050409-2